If in a triangle ABC a Cos*2(c/2)+cos*2(A/2)=3b/2 then the sides a,b,c are

Grade:12th pass

1 Answers

ALL IN ONE
6 years ago
we know cosC=2cos^2(C/2)-1
 
so, a(1+cosC)/2+c(1+cosA)/2=3b/2
 
a+acosC+c+ccosA=3b
 
a+c+(acosC+ccosA)=3b
 
Using By projection formula: (acosC+ccosA)=b
 
a+c+b=3b
 
a+c=2b
This shows that a,b,c are in A.P.
thank you. my name : ashu mehta (bupp), (sirsa)